I was wondering if you would be able to assist me with this problem...
y=x^2-2x-9

Answer by edjones(8007)

x^2-2x-9=0
x^2-2x =9
x^2-2x+1=9+1 comleting the square by dividing the middle term by 2 and squaring it and last term.
(x-1)^2=10
x-1=+-sqrt(10) sqrt of each side.
x=1+-sqrt(10)
.
Ed